OpenStudy (anonymous):

Simplify the expression: 20 / 4*9

OpenStudy (anonymous):

@monoremix

OpenStudy (anonymous):

20/36 5/9 <- answer

OpenStudy (anonymous):

it tells me its wrong

OpenStudy (anonymous):

now it gave me a new problem 21 divided by 3 times 9

OpenStudy (anonymous):

21/3=7*9=63 63 is the answer

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

welcome..